OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

hartnn (hartnn):

put n=3 in f(n) = f(n - 1) + f(n - 2), what u get ?

OpenStudy (anonymous):

f(3) = f(2)+f(1) ?

OpenStudy (chihiroasleaf):

yes.., it is.. :)

OpenStudy (gorv):

f(3)=f(2)+f(1)=1 f(4)=f(3)+f(2)=2 f(5)=f(4)+f(3)=3

OpenStudy (anonymous):

I don't get which answer it is though

OpenStudy (gorv):

in Question put n=3 , 4, 5

OpenStudy (gorv):

u will get ur ans

OpenStudy (loser66):

still not get?

OpenStudy (gorv):

f(n) = f(n - 1) + f(n - 2) put n=3

OpenStudy (gorv):

tell me wat u are getting

OpenStudy (anonymous):

no i get it but i just don''t get if its a b c or d

OpenStudy (gorv):

number will given in place of a b c d

OpenStudy (anonymous):

ok thanks how do i do this one? Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______